OpenStudy (anonymous):

sinx=1/3 and x lies in quad. 2, find exact value of sin(x+pi/6) THANKS!

OpenStudy (jdoe0001):

\(\bf sin(x) = \cfrac{1}{3} \implies \cfrac{\textit{opposite}}{\textit{hypotenuse}} \implies\cfrac{b}{c}\\ c^2 = a^2 + b^2 \implies \sqrt{c^2-b^2} = a\\ \textit{keep in mind that}\\ \color{blue}{sin(\alpha-\beta) = sin(\alpha)cos(\beta)-cos(\alpha)sin(\beta)}\)

OpenStudy (jdoe0001):

also recall that is "x" is in the 2nd Quadrant, "x" is negative, "y" is positive so the sine will be positive, as you can see it's but the cosine will be negative

OpenStudy (jdoe0001):

well, I'd say angle "x" is in the 2nd Quadrant and cosine there is negative

OpenStudy (jdoe0001):

lemme word it better :) also recall that is "GIVEN ANGLE" is in the 2nd Quadrant, "x" is negative, "y" is positive there so the sine will be positive, as you can see it's but the cosine will be negative

OpenStudy (anonymous):

Thanks but I'm still a bit confused as to how to find x

OpenStudy (jdoe0001):

well, you're not required to find the angle "x" you're just required to find the exact value of the sine function of the angle "x + pi/6"

OpenStudy (jdoe0001):

and you'd do it by using the coordinate values for "x"

OpenStudy (jdoe0001):

if we know that \(\bf sin(x) = \cfrac{1}{3} \implies \cfrac{\textit{opposite}}{\textit{hypotenuse}} \implies\cfrac{b}{c}\\ c^2 = a^2 + b^2 \implies \sqrt{c^2-b^2} = a\) what value would you get for "a"?

OpenStudy (anonymous):

A= radical 8

OpenStudy (jdoe0001):

\(\bf a = \pm \sqrt{8} \implies a = \pm 2\sqrt{2}\) keep in mind, "ANGLE x" is in the 2nd Quadrant, so it's cosine will be negative, thus \(\bf a = -2\sqrt{2}\)

OpenStudy (jdoe0001):

so now that we know that, what would be the cosine of "x" what's is cos(x)?

OpenStudy (anonymous):

.999?

OpenStudy (jdoe0001):

how did you get that?

OpenStudy (anonymous):

I put cos(-2rad2) into calculator

OpenStudy (jdoe0001):

hmmm, well, we didn't find the cos(x) btw we only found the side "a" for the angle "x"

OpenStudy (jdoe0001):

cosine of the angle will be \(\bf cos(x) = \cfrac{\textit{adjacent side}}{\textit{hypotenuse}}\implies\cfrac{a}{c}\)

OpenStudy (anonymous):

Ohhh ok

OpenStudy (jdoe0001):

so, what's the hypotenuse for "x"?

OpenStudy (anonymous):

Sry im notat home right now soim responding from my phone

OpenStudy (jdoe0001):

ok

OpenStudy (anonymous):

wow thanks so much! i finally got it and got the right answer :D